rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2972 — FS#6932 — vps-20-n5 et vps-21-n5

Attached to Project— VPS
Maintenance
Infrastructure
CLOSED
100%
We are going to update the switches managing the offers VPS and PCI.
The upgrade will be done on 2 ways independently one after the other. There is no expected interruption.

We planned the update in the night of Wednesday 27th June to Thursday 28th June at 00:00 (CEST).
Date:  Thursday, 26 July 2012, 05:16AM
Reason for closing:  Done
Comment by OVH - Thursday, 28 June 2012, 13:29PM

We start with the switch vps-21-n5.


Comment by OVH - Thursday, 28 June 2012, 13:30PM

The update was done with success. All the trafic is on vps-20-n5.

The switch vps-21-n5 is now updated and operational.

We switch the vps-20-n5 on vps-21-n5.


Comment by OVH - Thursday, 28 June 2012, 13:33PM

All the trafic is switched on vps-21-n5.

We are updating the switch vps-20-n5 with its Fex.


Comment by OVH - Thursday, 28 June 2012, 13:33PM

The switch vps-20-n5 is updated.

Everything is operational.


Comment by OVH - Tuesday, 24 July 2012, 16:37PM

We detect a bug in the version of the 2 switchs.

2012 Jul 24 07:10:40 vps-20-n5 %ETHPORT-5-IF_SEQ_ERROR: Error ("sequence timeout") communicating with MTS_SAP_STP for opcode MTS_OPC_ETHPM_PORT_LOGICAL_CLEANUP (RID_PORT: Ethernet107/1/18) - Ethpm didn't receive a Link UP from the driver on time or had an
internal error, check Port Client and Ethpm logs and debug

We will proceed to their update during the night of the 25th of July till the 26th at 00:00.
The infrastructure as it is redundant, no interruption of traffic is expected.


Comment by OVH - Thursday, 26 July 2012, 02:23AM

We are switching traffic on vps-21-n5.


Comment by OVH - Thursday, 26 July 2012, 02:23AM

Traffic is now switched.

We will update vps-20-n5.


Comment by OVH - Thursday, 26 July 2012, 03:18AM

The switch is updated.

We are switching the traffic on the updated switch vps-20-n5.


Comment by OVH - Thursday, 26 July 2012, 03:19AM

The traffic has been switched.

We are updating vps-21-n5.


Comment by OVH - Thursday, 26 July 2012, 05:16AM

Both switches have been updated and are running.